hello I need help woth the following questions.. please show your work
1.
2.

3.

1
2

3
Harold Corporation, a manufacturer of electronics and communications systems, uses a service department charge system to charge profit centers with Computing and Communications Services (CCS) service department costs. The following table identifes an abbreviated list of service categories and activity bases used by the CCS department. The table also includes some assumed cost and activity base quantity information for each service for April. CCs Service Budgeted Activity Base Quantity Category Activity Base Budgeted Cost Help desk Number of calls s88,0400 3.100 Number of devices monitored Network center 474,525 8,550 Electronic mal Number of user accounts 64.500 6,450 128.000 s.000 Local volce support Number of phone extensions One of the profit centers for Harold Corporation is the Communication Systems (COMM) sector. Assume the following information for the COMM sector: The sector has 1,000 employees, of whom 40% are office employees All the office employees have a phone, and 90% of them have a computer on the network 90 percent of the employees with a computer also have an e-mail account. The average number of help desk calls for April was 1 call per individusl with a computer. There are 300 addtional printers, servers, and peripherals on the network beyond the personal computers a. Determine the service charge rate for the four CCS service categories for April Round your answers to two decimal places Service Charge Rate ccs Service Categery 284 Per call Help desk 55.5 Per device monitored Network center Per e-mail account Electronic mail 10 16 Per phone extension Local voice support b. Determine the charges to the COMM sector for the four CCS service categoeries for April April charges to the COMM sector 9.368.6 Help desk charge 115,200 X Network center charge Electronic mal charge 17.034 X 6.400 Local volce support charge
Decision on Transfer Pricing Materials used by the Truck Division of Armatrong Motors are currently purchased from outside suppliers at a cost of $290 per unit. However, the same materials are available from the Components Division. The Components Division has unused capacity and can produce the materials needed by the Truck Division at a variable cost of $241 per unit a. Ifa transfer price of $264 per unit is established and 32,000 units of materials are transferred, with no reduction in the Components Dvision's current sales, how much would Armstreng Motors' total income from operations increase b. How much would the Truck Division's income from operations increase" How much would the Compenents Division's income from operations increase
Rate of Return on Investment Commodore Entertainment recently reported sector income from operations, revenue, and invested assets as follows: Income from Invested Operations Revenuer Assets Media Networks $977,000 $167,067 $781,600 Parks and Resorts 56.810 418.600 598.000 12.312 324,000 Studio Entertainment 291.600 Consumer Products 100,345 423,000 235,000 Use the DuPont formula to determine the rate of return on investment for the four Commodore Entertainment sectors. Round Proft Margin, ROI to one decimal place and Investment Turnover to two decimal places Profit Margin Investment Turnover ROI 21.37 Meda Networks 17.1 Parks and Resorts 13.57 95 x% Studio Entertainment 4 38 X% Consumer Products 23.72 18 42.7
W CH 24 Harold Corporation, a manufacturer of electronics and communications systems, uses a service department charge system to charge profit centers with Computing and Communications Services (CCS) service department costs. The following table identifies an abbreviated list of service categories and activity bases used by the ccs department. The table also includes some assumed cost and activity base quantity information for each service for April. Budgeted Activity Base Quantity CCS Service Budgeted Cost Activity Base Category 3,100 $88,040 Number of calls Help desk 8,550 474,525 Number of devices monitored Network center 6,450 64,500 Number of user accounts Electronic mail 8,000 128,000 Local voice support Number of phone extensions One of the profit centers for Harold Corporation is the Communication Systems (COMM) sector. Assume the following information for the COMM sector: The sector has 1,000 employees, of whom 40% are office employees. All the office employees have a phone, and 90 % of them have a computer on the network 90 percent of the employees witha computer also have an e-mail account The average number of help desk calls for April was 1 call per individual with a computer. There are printers, servers, and peripherals on the network beyond the personal computers. 300 addition a. Determine the service charge rate for the four CCS service categories for April. Round your answers to two decimal places. Service Charge Rate CCS Service Category 28.4 Per call Help desk Per device monitored 55.5 Network center Per e-mail account 10 Electronic mail Per phone extension 16 Local voice support b. Determine the charges to the COMM sector for the four CCS service categories for April. April charges to the COMM sector: X 9,368.6 Help desk charge 115,200 X Network center charge 17,034 Electronic mail charge 6,400 Local voice support charge Novt
Rate of Return on Investment Commodore Entertainment recently reported sector income from operations, revenue, and invested assets as follows: Invested Income from Assets Revenue Operations $781,600 $977,000 $167,067 Media Networks 598,000 418,600 56,810 Parks and Resorts 324,000 291,600 12,312 Studio Entertainment 235,000 100,345 423,000 Consumer Products Use the DuPont formula to determine the rate of return on investment for the four Commodore Entertainment sectors. Round Profit Margin, ROI to one decimal place and Investment Turnover to two decimal places. ROI Profit Margin Investment Turnover 17.1 21.37 % X 8 Media Networks 7 X 13.57 % 95 X % Parks and Resorts X % 38 4 % Studio Entertainment 42.7 1.8 23.72 % Consumer Products Feadback Y Check Mhy Vor Income from operations divided by salos equals profit margin. Sales divided by invested assets equals investment turnover. Multiply these two percentages for the rate of return Leaming Obiective 4
Decision on Transfer Pricing Materials used by the Truck Division of Armstrong Motors are currently purchased from outside suppliers at a cost of $290 per unit. However the same materials are available from the Components Division. The Components Division has unused capacity and can produce the materials needed by the Truck Division at a variable cost of $241 per unit. a. If a transfer price of $264 per unit is established and 32,000 units of materials are transferred, with no reduction in the Components Division's current sales, how much would Armstrong Motors' total come from operations increase? b. How much would the Truck Division's income from operations increase? c. How much would the Components Division's income from operations increase?
Answer #1

1. Harold Corporation

Help Desk Charge = Number of Employees * % of Office Employees * % of Computer Network * Help desk Service Charge Rate

Help Desk Charge = 1000 * 40% * 90% * $28.40

Help Desk Charge = $10224

Network Center Charge = ((Number of Employees * % of Office Employees * % of Computer Network)+300) * Network Center Service Charge Rate

Network Center Charge = ((1000 * 40% * 90%)+300) * $55.50

Network Center Charge = $36630

Electronic Main Charge = (Number of Employees * % of Office Employees * % of Computer Network * % of Email Account) * Electronic Mail Service Charge Rate

Electronic Main Charge = (1000 * 40% * 90% * 90%) * 10

Electronic Main Charge = $3240

2. Commodore Entertainment

Investment Turnover = Revenue / Invested Assets

Media Networks = 781600 / 977000 = 0.80

parks and Resorts = 418600 / 598000 = 0.70

Studio Entertainment = 291600 / 324000 = 0.90

ROI = Income From Operations / Invested Assets

parks and Resorts = 56810 / 598000 = 9.5%

Studio Entertainment = 12312 / 324000 = 3.8%

3.

a. Increase in Income of Operations = Units Transferred * (Purchase Price - Variable Cost)

Increase in Income of Operations = 32000 * (290 - 241)

Increase in Income of Operations = $1568000

b Truck Division Increase in Income = Units Transferred * (Purchase Price - transfer price)

Truck Division Increase in Income = 32000 * (290 - 264)

Truck Division Increase in Income = $832000

c Components Division Increase in Income = Units Transferred * (transfer Price - cost)

Components Division Increase in Income = 32000 * (264 - 241)

Components Division Increase in Income = $736000